OpenStudy (anonymous):

Find two constants a and b such that the test item "Simplify a(x+2y)-b(2x-y)" will have an answer of 3x+9y. what are the constants a and b?

myininaya (myininaya):

you have a(x+2y)-b(2x-y) what do you get when you multiply a(x+2y)?

myininaya (myininaya):

i will do this part for you, but you do the next a(x+2y)=a*x+a*2y=ax+2ay

myininaya (myininaya):

now multiply b(2x-y)

OpenStudy (anonymous):

2bx-by

OpenStudy (anonymous):

sorry correction -2bx+by

myininaya (myininaya):

ok so you decided to multiply -b(2x-y) which is even better! gj

myininaya (myininaya):

so we have ax+2ay-2bx+by right?

OpenStudy (anonymous):

right

myininaya (myininaya):

ok now we need need to get to the answer 3x+9y so we need to put our x's together and put our y's together so i'm just gonna move some things around k? ax-2bx+2ay+by is this okay with you?

OpenStudy (anonymous):

I follow

myininaya (myininaya):

ok now can i do this x(a-2b)+y(2a+b) ?

OpenStudy (anonymous):

dosent that equal 3x+9y

myininaya (myininaya):

yes we want this to equal 3x+9y so we want the 3 to be a-2b and we want the 9 to be 2a+b so we have two equations: a-2b=3 and 2a+b=9 do you know how to solve a system of equations?

OpenStudy (anonymous):

Matrix elimination

OpenStudy (anonymous):

is it 4.2 and 0,6

myininaya (myininaya):

ok yes i think thats what it is called what if we multiply the second equation by 2: 2(2a+b=9) so the 2nd equation becomes 4a+2b=18 so we have a -2b= 3 4a+2b=18 so what if we add the equations togehter 5a+0=21 5a=21 a=21/5 so we can find b now pluggin a into one of the above equations

myininaya (myininaya):

gj chuddley :)

OpenStudy (anonymous):

thank you very much and you were great help as well myininaya

myininaya (myininaya):

:)